40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 6 20:55:08 2013, rana, Summary, ASC, MC WFS Limits set based on 40 days of trends mcwrfs_trend.png
    Reply  Sun Oct 6 22:27:35 2013, rana, Frogs, Electronics, MC3 LL sensor cable was loose MC3_LL.png
Message ID: 9208     Entry time: Sun Oct 6 22:27:35 2013     In reply to: 9207
Author: rana 
Type: Frogs 
Category: Electronics 
Subject: MC3 LL sensor cable was loose 

I noticed that the MC3 LL sensor was apparently dead according to its suspension screen. Since it was only the fast ADC channel and not the SLOW PDmon, I could tell that it was just in the ADC cabling. I pushed in a few of the MC3 sensor cables on the front and back of the PD whitening board and it came back OK. According to this trend of the past 40 days and 40 nights, it started slipping on this past Wednesday morning.

Was anyone walking near MC2 or the suspension electronics racks before noon on Wednesday (Oct. 2nd)?

Attachment 1: MC3_LL.png  10 kB  | Hide | Hide all
ELOG V3.1.3-